# Traffic-R1: Reinforced LLMs Bring Human-Like Reasoning to Traffic Signal Control Systems

### v1

@@ 5.3. OOD Incident Dealing Performance (RQ2)
Table 3. OOD Incident Dealing Performance Evaluation. “-” indicates the method is entirely inadequate for the task.
@@
DeepSeek-R1 672B 0.93 223.19 10.14
Traffic-R1 3B 0.85 215.58 7.98
5.4. Discussion of RL-based Finetuning (RQ3)
@@
We proposes a new RL-based two-stage finetuning paradigm for training LLMs instead of traditional instruction finetuning in TSC tasks. To further validate its advantages, we conduct detailed comparisons and discussion in this section. We adopt the LLMLight (Lai et al., 2023) as a representative framework of instruction finetuning. For a fair comparison, we upgrade the base model of LLMLight from Qwen2 to Qwen2.5, matching the base model of Traffic-R1. We follow the same pipeline as Traffic-R1, using traditional RL methods to search for optimal action trajectories in our open-world RL environment. Additionally, we use GPT-4 to generate explanations for trajectories and expert-collaborative QA samples to construct the instruction dataset. After training, we evaluate the models in two dimensions: 1. TSC performance: We assess the zero-shot TSC capabilities of the two fine-tuned LLMs and the base model, quantifying the results by setting the base model’s performance at 0.5. The performance of the other models is computed based on their average improvement in ATT and AWT across all datasets. 2. General capabilities: we evaluate reasoning, instruction following, and inner commonsense on public benchmarks (Team, 2024) to examine the models’ foundational capacities, these capacities further indicate the potential and stability of the finetuned model for real-world wild deployment. The results, shown in Figure 5, highlight two key findings regarding RL-based finetuning for LLMs in traffic systems:
Figure 5. Comparison results for models’ capacities.
  • From imitation to human-like reasoning. Traffic-R1 achieves significantly better zero-shot performance across various TSC tasks compared to the instruction finetuned model, which exhibits unsatisfactory performance under the same zero-shot settings. The performance gap, despite identical training data and environments, reveals the limitations of instruction finetuning: models tend to memorize and imitate synthetic training data unless provided with a large, high-quality, and comprehensive dataset. In contrast, our RL-based finetuning encourages models to develop internal reasoning by promoting self-exploration and iteration during training, ensuring superior zero-shot TSC performance.
  • Complete general capabilities. Without a comprehensive dataset and tailored training settings, instruction finetuning often causes LLMs to lose general capabilities while adapting to training samples. This issue is particularly pronounced in traffic tasks, which differs significantly from general LLM tasks. The model trained under the LLMLight framework shows a substantial reduction in general capabilities and performs worse in other scenarios compared to base model. However, our RL-based framework mitigates catastrophic forgetting through self-rollout samples and KL divergence guidance. This approach constrains the base model to evolve within a controlled parameter space, avoiding simple memorization and promoting deeper, more efficient optimization for better comprehensive performances.
5.5. Ablation Study (RQ4)
@@
Figure 6. Ablation study on conventional and OOD tasks.
Figure 7. Online and offline dispatch framework
6. Real-world deployment
Pipeline. Traffic-R1 is deployed on a major real-world traffic platform in a large Chinese city (name withheld for privacy), managing 10 key intersections and serving over 55,000 drivers daily. These intersections are primary junctions responsible for four directions within the road network and are interconnected by distances of approximately 600 to 800 meters. The road network is located in a commercial area of the city, characterized by heavy and variable traffic flow loads during different peak hours. The peak daily vehicle flow at one intersection can reach up to 21,000, indicating a significant traffic load. Unlike traditional simulated environments, these real-world main intersections have more lanes in each direction. To simplify the complexity of inputs and reasoning, we merge the lanes with the same flow direction. Additionally, a passenger phase is strictly added to the phase list at these intersections during morning and evening commute hours. This adjustment disrupts the standard phases and requires a robust TSC capacity from the model to effectively manage the increased complexity.
Integrating Traffic-R1 into a real-world TSC system presents several challenges, with the first being the accurate sensing of real-time traffic conditions and their conversion into formats readable by the LLM. To address this, we developed an online traffic sensing system that utilizes a finetuned Grounding DINO model for traffic object recognition, along with millimeter-wave radar to measure the distance of objects from intersections. As shown in Figure 8, a road camera mounted on the traffic light provides clear 2D visual information of vehicles in each lane, while the millimeter-wave radar complements this by measuring the distance between vehicles and the intersection.
Figure 8. In-suit traffic sensing devices on road intersections
While the sensing system provides continuous traffic observation for each intersection with the same format of simulation environment, deploying the model in the real world cannot rely on the straightforward, fully online mechanisms used in simulations due to mandatory government oversight for safety and security. As illustrated in Figure 7, we designed an integrated online and offline dispatch framework to meet these practical requirements, enabling Traffic-R1 to function as both a traffic marshal and a traffic plan designer. The online dispatch pipeline handles real-time incident response, while the offline dispatch pipeline manages routine TSC operations. In the offline pipeline, traffic observations are abstracted into spatio-temporal traffic flow representations for each intersection. These flows are analysis by a LGBM algorithm with spatio-temporal embeddings for predicting further trends. Specifically, we use 30 days of historical data to forecast traffic patterns for the next day. Traffic-R1 then generates a preliminary action plan based on these predictions. This plan undergoes human evaluation and revision by relevant authorities before being implemented.
Post-launch Performance. We evaluate the post-launch performance of Traffic-R1 through a parallel experiment comparing our pipeline with original pipeline. The original pipeline relies on human-labor with for action plan design and incident dealing. We deploy the two pipelines to perform traffic control on an alternating weekly basis for a total of 6 weeks, and record the daily average, maximum queue length of busy hours (4pm to 7:30 pm) and human working hours of all intersections in the network, as presented in Table 4. To avoid interference, data collected during weekend and holidays are discarded. Results show the Traffic-R1 significantly enhances the real-world traffic control system by efficient human-like reasoning and decisions, which greatly exceed the working efficiency of human experts and save human labor with stable performance.
Table 4. Real-world parallel test results spanning 6 weeks.
@@ 6. Real-world deployment
#Improvement 9.3% 4.4% $\sim$75%
7. Conclusion and Future Work
In this paper, we present Traffic-R1, a reinforced LLM designed for real-world TSC systems, capable of human-like reasoning. We propose a novel two-stage reinforcement learning strategy and a network communication framework to develop the LLM into a foundational model for traffic control that functions like a human traffic agent. Evaluations on conventional TSC and traffic incident handling demonstrate its superiority over existing methods, while real-world experiments further indicate its value for industrial deployment. Future work can focus on exploring reinforced vision-language models to enable direct reasoning and decision-making based on road vision information, enhancing efficiency and deployment convenience. Additionally, efforts can aim to mitigate hallucination-related errors caused by LLMs within the communication process in the traffic network through more robust network designs or model-level solutions.
@@ Appendix A Appendix
A.1. Input Template for Rollout
Input Templates for LLM Rollout System: You are a helpful traffic control agent.
TSC Description: The crossroad connects two roads: north-south and east-west, with the traffic light at their intersection. Each road is divided into two sections (e.g., north and south for the north-south road) and each section has two lanes: a through lane and a left-turn lane…
Traffic Observation: Signal: ETWT
Allowed lanes: Eastern and western through lanes
- Early queued: 2 (East), 1 (West), 3 (Total)
- Segment 1: 0 (East), 0 (West), 0 (Total)
- Segment 2: 1 (East), 0 (West), 1 (Total)
@@ A.1. Input Template for Rollout
A.2. Traffic Incident Data
Examples for testing traffic incidents ID 1: At this intersection, a traffic accident in the eastbound lane is causing significant congestion. Response Action: ETWT
ID 2: At this intersection, a pedestrian was struck in the northbound crosswalk. Response Action: ETWT/NLSL
ID 3: Report from the nearby intersection to the north: Heavy southbound traffic is approaching. Response Action: NTST/NLSL
ID 4: At this intersection, a school bus is stopped in the eastbound lane, loading students. Response Action: NTST
ID 5: At this intersection, a group of pedestrians is blocking the westbound crosswalk. Response Action: NTST
ID 6: Report from the nearby intersection to the east: A fire hydrant crew is slowing westbound traffic. Response Action: NTST/ELWL
ID 7: At this intersection, vehicles spun out in the westbound lane. Response Action: NTST
ID 8: At this intersection, for the XXX event, a marathon is passing through the eastbound lane. Response Action: ETWT
ID 9: At this intersection, a road rage incident was reported in the northbound lane due to a heavy traffic jam. Response Action: NTST
ID 10: At this intersection, it is 5:30 PM, the nearby western school is scheduled to dismiss, leading to increased vehicular traffic for student pick-up. Response Action: ETWT
Table 5. Performance comparison on conventional traffic signal control tasks (the smaller the better). The best results are in bold and second-best results are underlined.
@@ A.3. Introduction of baselines
    • FixedTime (Koonce et al., 2008): A policy that assigns a fixed cycle length with predefined phase splits across all phases.
    • MaxPressure (Varaiya, 2013): A control strategy that selects the phase with the highest pressure to optimize traffic flow.
@@
    • LLMLight-7B (Lai et al., 2023): A SOTA LLM-based TSC method that employs the Advanced-CoLight framework to interact and generate action policies for each dataset. It utilizes GPT-4 to generate explanations for each action, which, along with Advanced-CoLight generated actions, are used for instruction finetuning to enable the LLM base model to emulate TSC capabilities.
    • MPLight (Chen et al., 2020): A method based on the FRAP model that uses pressure as both observation and reward to optimize TSC.
    • AttendLight (Oroojlooy et al., 2020): A method that employs attention mechanism to construct phase features and predict its transition probabilities.
    • CoLight (Wei et al., 2019): A method that uses a graph attention network to represent inter-intersection communication within a RL framework.
    • Efficient-CoLight (Wu et al., 2021): An enhanced version of the CoLight model that incorporates efficient pressure as an observation to improve decision-making in TSC.
    • Advanced-CoLight (Zhang et al., 2022): A SOTA RL-based method that enhances CoLight by integrating efficient pressure and advanced traffic state features, such as effective running vehicles, to optimize decision-making capabilities.
@@
    • CoLLMLight-8B (Yuan et al., 2025): An enhanced version of LLMLight (Lai et al., 2023) that incorporates neighbor information into consideration. The model is trained using an instruction fine-tuning framework on a synthetic dataset and demonstrates improved zero-shot performance compared to LLMLight.
@@
Responses of Traffic-R1 in different scenarios Case 1: ¡think¿Based on the current state of the intersection, Signal: ETWT (Eastern and Western through lanes) is the one that most significantly improves traffic conditions during the next phase. This is because this lane has 18 early queued vehicles and 13 approaching vehicles in Segment 3, indicating a significant number of vehicles waiting to pass through and incoming huge traffic pressure. Approving this signal will allow these vehicles to proceed, easing the traffic flow.¡/think¿
Case 2: ¡think¿Based on the current state of the intersection, the signal phase that will most significantly improve traffic conditions during the next phase is NTST (Northern and southern through lanes). This is because only the southern through lane has approaching vehicles in Segment 1, which indicates that there are vehicles waiting to enter the intersection. By allowing vehicles in the northern and southern through lanes to flow, this signal phase will help avoid the queue of approaching vehicles and improve overall traffic flow.¡/think¿
Case 3: ¡think¿Based on the current state of the intersection, the signal phase that will most significantly improve traffic conditions during the next phase is ETWT (Eastern and Western through lanes). Although the northern through lane has 20 early queued vehicles in Segment 2, an ambulance is currently approaching from the east, heading west. Prioritizing ETWT is critical to ensure the ambulance passes through the intersection without delay, maintaining emergency response efficiency. This signal will also alleviate some traffic pressure by allowing eastern and western vehicles to proceed.¡/think¿
Case 4: ¡think¿Based on the current state of the intersection, the signal phase that will most significantly improve traffic conditions during the next phase is NTST (Northern and Southern pedestrian crossing). This is because a large number of students are waiting to cross from the northern to southern direction, indicating significant pedestrian demand. Activating this signal will allow these students to cross safely, reducing congestion at the crosswalk.¡/think¿
@@
Case 5: ¡think¿Based on the current state of the intersection and the urgent message of an emergency fire truck approaching from the west, the signal phase that will most significantly improve traffic conditions during the next phase is ETWT (Eastern and Western through lanes). This is because both ETWT and NTST have comparable queues, with approximately 15 early queued vehicles each in Segment 2, but prioritizing ETWT ensures the fire truck’s swift passage. Activating this signal will clear the western lane, facilitating emergency access while also easing general traffic flow.¡/think¿
A.4. More Experiment Result
@@
A.5. Reasoning Cases of Traffic-R1

### v2

@@ 5.3 OOD Incident Dealing (RQ2)
Table 3: OOD Incident Dealing Evaluation. “-” indicates the method is entirely inadequate for the task.
@@
[…]
We split OOD tasks into two types: local intersection incidents (single-intersection events, e.g., an local accident) and network-wide incidents (events affecting multiple intersections, e.g., emergency-vehicle routing). For local incidents we use Emergency Action Accuracy (EAA). For network-wide incidents we use Average Emergency Travel Time (AETT) and Average Emergency Waiting Time (AEWT) adapted from Wang et al. (2024). Results are reported in Table 3. The findings highlight two key strengths of Traffic-R1:
  • Stable OOD generalization. Traffic-R1 maintains consistent performance across diverse OOD scenarios and outperforms larger general LLMs and LLMLight by over 30% across metrics, indicating it applies traffic knowledge instead of merely imitating training examples.
  • Lightweight and efficient. At 3B parameters, Traffic-R1 matches or exceeds traditional baselines and advanced LLMs while using roughly 1% of their parameters and much lower deployment cost. It performs particularly well on network-wide incidents, showing that the asynchronous communication design enables coordinated multi-intersection control.
5.4 Discussion of RL-based Finetuning (RQ3)
@@
To validate our proposed RL-based two-stage finetuning paradigm, we compare it against traditional instruction finetuning, represented by the LLMLight framework Lai et al. (2023). For a fair comparison, both methods use the same Qwen2.5 base model. We built the instruction dataset for LLMLight using a similar pipeline: RL-generated action trajectories, GPT-4 explanations, and expert-collaborative QA samples. After training, we evaluate the models in two dimensions: 1. TSC performance: we assess zero-shot capabilities, quantifying results based on the average improvement in ATT and AWT across all datasets, with Qwen2.5-3B’s performance normalized to 0.5. 2. General capabilities: we use public benchmarks Team (2024) to evaluate reasoning, instruction following, and commonsense to gauge the models’ stability and potential for real-world deployment. The results in Figure 5, highlight two key findings regarding RL finetuning for LLMs in traffic tasks:
Figure 5: Comparison results for models’ capacities.
  • From imitation to reasoning. Traffic-R1 significantly outperforms the instruction-finetuned model on zero-shot TSC tasks. This performance gap, despite identical training data, shows that instruction finetuning encourages imitation of synthetic data. In contrast, our RL-based paradigm fosters internal reasoning through self-exploration and iteration.
  • Complete general capabilities. Instruction finetuning often causes "catastrophic forgetting," degrading LLM’s general abilities. The SFT model exemplifies this, performing worse than the base model on general benchmarks. Traffic-R1 mitigates this issue using self-rollout samples and KL divergence penalty. This approach constrains the policy update to a controlled space, promoting deeper optimization over memorization and preserving the model’s comprehensive capabilities.
5.5 Ablation Study (RQ4)
@@
Figure 6: Ablation results on TSC and OOD tasks.
6 Real-world deployment
Figure 7: In-suit traffic sensing devices for Traffic-R1
As shown in Figure 7, Traffic-R1 is deployed on a online traffic platform in a Chinese city (name withheld for privacy), controlling 10 key intersections and serving over 55,000 drivers daily. A/B tests for 6 weeks (Table 6) show significant improvements in real-world traffic control—more efficient decision-making, superior performance compared with human experts, and consistent operation with reduced labor. We place details and discussion in Appendix A.6 due to space limit.
Table 4: A/B test results spanning 6 weeks.
@@ 6 Real-world deployment
#Improve 9.3% 4.4% $\sim$75%
7 Conclusion and Future Work
In this paper, we introduce a two-stage reinforcement learning strategy and a network communication framework to convert the LLM into a foundational traffic-control model that operates like a human traffic agent. Evaluation on standard TSC benchmarks and traffic-incident handling shows improvement over prior methods, and real-world tests demonstrate value for industrial deployment. Future work includes reinforced VLMs for direct road-vision reasoning without reliance on structured textual traffic data, enhancing efficiency and deployment convenience.
8 Limitations
While Traffic-R1 demonstrates strong performance in both standard and out-of-distribution traffic control tasks with impressive zero-shot generalization, it has several limitations. First, the base model, training, and inference are highly reliant on structured data in a single textual modality: traffic observations and incident information must be converted into structured text before being input to the model. This requirement causes inconvenience and potential information loss in dense real-world deployments, and we will address it in future work by developing reinforced vision-language models and multimodal input pipelines. Second, due to limitations in deployment scale and practical constraints imposed by traffic management agencies, we cannot provide a detailed analysis of LLM hallucination-related instability across large-scale traffic scenarios or assess drivers’ feedback to the agent-based traffic network. Access restrictions, privacy requirements, safety constraints, and the cost and complexity of long-term field trials prevented extended real-world evaluations and systematic collection of driver response data.
@@ Appendix A Appendix
A.1 Input Template for Traffic-R1
Input Templates for LLM Rollout System: You are a helpful traffic control agent. Task Description: The crossroad connects two roads: north-south and east-west, with the traffic light at their intersection. Each road is divided into two sections (e.g., north and south for the north-south road) and each section has two lanes: a through lane and a left-turn lane… Structured Traffic Observation:
Signal: ETWT
Allowed lanes: Eastern and western through lanes - Early queued: 2 (East), 1 (West), 3 (Total) - Segment 1: 0 (East), 0 (West), 0 (Total) - Segment 2: 1 (East), 0 (West), 1 (Total) - Segment 3: 1 (East), 2 (West), 3 (Total) Signal: ELWL Allowed lanes: Eastern and western left lanes - Early queued: 0 (East), 1 (West), 1 (Total) - Segment 1: 0 (East), 0 (West), 0 (Total) - Segment 2: 3 (East), 0 (West), 3 (Total) - Segment 3: 1 (East), 0 (West), 1 (Total) Signal: NTST Allowed lanes: North and south through lanes - Early queued: 0 (North), 0 (South), 0 (Total) - Segment 1: 0 (North), 0 (South), 0 (Total) - Segment 2: 1 (North), 0 (South), 1 (Total) - Segment 3: 1 (East), 1 (West), 2 (Total) Signal: NLSL Allowed lanes: North and south left lanes - Early queued: 1 (North), 0 (South), 1 (Total) - Segment 1: 0 (North), 0 (South), 0 (Total) - Segment 2: 0 (North), 0 (South),0 (Total) - Segment 3: 1 (East), 0 (West), 1 (Total) Incident Information (optional):
(at training stage we do not use this part)
@@ A.1 Input Template for Traffic-R1
[…]
@@ A.2 Introduction of baselines
    • FixedTime Koonce et al. (2008): A policy that assigns a fixed cycle length with predefined phase splits across all phases.
    • MaxPressure Varaiya (2013): A control strategy that selects the phase with the highest pressure to optimize traffic flow.
@@
    • LLMLight-7B Lai et al. (2023): A SOTA LLM-based TSC method that employs the Advanced-CoLight framework to interact and generate action policies for each dataset. It utilizes GPT-4 to generate explanations for each action, which, along with Advanced-CoLight generated actions, are used for instruction finetuning to enable the LLM base model to emulate TSC capabilities.
    • MPLight Chen et al. (2020): A method based on the FRAP model that uses pressure as both observation and reward to optimize TSC.
    • AttendLight Oroojlooy et al. (2020): A method that employs attention mechanism to construct phase features and predict its transition probabilities.
    • CoLight Wei et al. (2019): A method that uses a graph attention network to represent inter-intersection communication within a RL framework.
    • Efficient-CoLight Wu et al. (2021): An enhanced version of the CoLight model that incorporates efficient pressure as an observation to improve decision-making in TSC.
    • Advanced-CoLight Zhang et al. (2022): A SOTA RL-based method that enhances CoLight by integrating efficient pressure and advanced traffic state features, such as effective running vehicles, to optimize decision-making capabilities.
@@
    • CoLLMLight-8B Yuan et al. (2025): An enhanced version of LLMLight Lai et al. (2023) that incorporates neighbor information into consideration. The model is trained using an instruction fine-tuning framework on a synthetic dataset and demonstrates improved zero-shot performance compared to LLMLight.
@@
Responses of Traffic-R1 in different scenarios Case 1: <think>Based on the current state of the intersection, Signal: ETWT (Eastern and Western through lanes) is the one that most significantly improves traffic conditions during the next phase. This is because this lane has 18 early queued vehicles and 13 approaching vehicles in Segment 3, indicating a significant number of vehicles waiting to pass through and incoming huge traffic pressure. Approving this signal will allow these vehicles to proceed, easing the traffic flow.</think> Case 2: <think>Based on the current state of the intersection, the signal phase that will most significantly improve traffic conditions during the next phase is NTST (Northern and southern through lanes). This is because only the southern through lane has approaching vehicles in Segment 1, which indicates that there are vehicles waiting to enter the intersection. By allowing vehicles in the northern and southern through lanes to flow, this signal phase will help avoid the queue of approaching vehicles and improve overall traffic flow.</think> Case 3: <think>Based on the current state of the intersection, the signal phase that will most significantly improve traffic conditions during the next phase is ETWT (Eastern and Western through lanes). Although the northern through lane has 20 early queued vehicles in Segment 2, an ambulance is currently approaching from the east, heading west. Prioritizing ETWT is critical to ensure the ambulance passes through the intersection without delay, maintaining emergency response efficiency. This signal will also alleviate some traffic pressure by allowing eastern and western vehicles to proceed.</think> Case 4: <think>Based on the current state of the intersection, the signal phase that will most significantly improve traffic conditions during the next phase is NTST (Northern and Southern pedestrian crossing). This is because a large number of students are waiting to cross from the northern to southern direction, indicating significant pedestrian demand. Activating this signal will allow these students to cross safely, reducing congestion at the crosswalk.</think> Case 5: <think>Based on the current state of the intersection and the urgent message of an emergency fire truck approaching from the west, the signal phase that will most significantly improve traffic conditions during the next phase is ETWT (Eastern and Western through lanes). This is because both ETWT and NTST have comparable queues, with approximately 15 early queued vehicles each in Segment 2, but prioritizing ETWT ensures the fire truck’s swift passage. Activating this signal will clear the western lane, facilitating emergency access while also easing general traffic flow.</think>
A.3 Traffic Incident Data
Examples for testing traffic incidents ID 1: At this intersection, a traffic accident in the eastbound lane is causing significant congestion. Response Action: ETWT ID 2: At this intersection, a pedestrian was struck in the northbound crosswalk. Response Action: ETWT/NLSL ID 3: Report from the nearby intersection to the north: Heavy southbound traffic is approaching. Response Action: NTST/NLSL ID 4: At this intersection, a school bus is stopped in the eastbound lane, loading students. Response Action: NTST ID 5: At this intersection, a group of pedestrians is blocking the westbound crosswalk. Response Action: NTST ID 6: Report from the nearby intersection to the east: A fire hydrant crew is slowing westbound traffic. Response Action: NTST/ELWL ID 7: At this intersection, vehicles spun out in the westbound lane. Response Action: NTST ID 8: At this intersection, for the XXX event, a marathon is passing through the eastbound lane. Response Action: ETWT ID 9: At this intersection, a road rage incident was reported in the northbound lane due to a heavy traffic jam. Response Action: NTST ID 10: At this intersection, it is 5:30 PM, the nearby western school is scheduled to dismiss, leading to increased vehicular traffic for student pick-up. Response Action: ETWT
A.4 More Experiment Result
@@
A.5 Reasoning Cases of Traffic-R1
@@
A.6 Deployment Details
Figure 8: Online and offline dispatch framework
A.6.1 Environment and System Adaption.
Our model, Traffic-R1, has been deployed on the official traffic management platform of a Tier-1 city in China, where it oversees 10 critical, interconnected intersections. This network serves over 55,000 vehicles daily and is situated in a central commercial district characterized by high-volume, dynamic traffic patterns. The intersections are configured as four-way junctions on arterial roads, with inter-intersection distances ranging from 600 to 800 meters. The traffic load is substantial, with peak daily throughput at a single intersection exceeding 21,000 vehicles.
To bridge the gap between simulation and real-world complexity, two key adaptations were implemented. First, to manage the increased number of lanes typical of urban arterial roads, we aggregated lanes with identical traffic movements into a single logical input, thereby reducing state-space complexity for the model. Second, the system incorporates a mandatory pedestrian-crossing phase during morning and evening peak hours. This regulatory requirement introduces additional constraints to the signal timing plan, demanding greater adaptive capabilities from our model to maintain traffic efficiency. Real-world deployment necessitates two critical components: a multi-modal sensing system and a safety-compliant dispatch framework.
To provide the model with real-time, LLM-compatible inputs, we developed a multi-modal sensing suite. As illustrated in Figure 7, roadside cameras provide 2D visual data, which is processed by a fine-tuned Grounding DINO model for vehicle detection and classification. This is augmented by millimeter-wave radar data that provides precise vehicle-to-intersection distance measurements, creating a comprehensive traffic state representation.
Direct, fully online model control is infeasible in public infrastructure due to regulatory and safety mandates. To address this, we designed a hybrid online-offline dispatch framework (Figure 8). The dispatch framework is executed on a centralized server equipped with an NVIDIA Tesla T40 24G GPU with an average inference latency of less than 2 seconds (437 output tokens in average). This centralized deployment is optimal for our predominantly offline dispatch strategy, allowing for robust data aggregation and computation. For routine traffic management, the system operates offline. Historical traffic data spanning 30 days is used to train a LightGBM model with spatio-temporal embeddings to forecast traffic flow for the subsequent day. Based on these forecasts, Traffic-R1 generates a preliminary signal timing plan. Crucially, this plan is submitted to human experts at the traffic management authority for review and final approval before implementation. The online pipeline is reserved for real-time incident response, allowing the model to make immediate adjustments under expert supervision. This dual-pipeline approach ensures both operational efficiency and compliance with public safety standards. Furthermore, while the current deployment leverages a center server, the model is designed for future scalability. The 3B-parameter Qwen architecture of Traffic-R1 can be quantized for efficient execution on low-power edge devices that lack dedicated GPUs.
Table 6: Real-world A/B test results spanning 6 weeks.
[…]
Manual 34.5 50.3 2+
Traffic-R1 31.3 48.1 0.5+
#Improve 9.3% 4.4% $\sim$75%
A.6.2 In-Situ Performance Evaluation
We conducted a six-week in-situ evaluation to quantify the performance of Traffic-R1. The experiment was structured as an A/B test, comparing our proposed system against the baseline, which consists of manual signal control by human operators. The two systems were deployed on an alternating weekly basis to mitigate temporal bias. Performance metrics, including average queue length, maximum queue length, and daily human operational hours, were recorded during weekday evening peak hours (16:00-19:30). Data from weekends and public holidays were excluded to ensure a consistent basis for comparison.
As detailed in Table 6, our system demonstrates marked improvements over the manual baseline. Specifically, Traffic-R1 reduced the average and maximum queue lengths by 9.3% and 4.4%, respectively. Furthermore, it achieved a substantial 75% reduction in the human labor required for daily operations. These results empirically validate the effectiveness and efficiency of our model in a real-world, dynamic traffic environment.
Algorithm 1 Asynchronous TSC Network with Symmetric Half-Step Communication
1:Input: Road network graph $G=(\mathcal{I},E)$, partitioned intersection sets $(\mathcal{G}_{1},\mathcal{G}_{2})$.
2:Initialize: Environment EnvEnv, Agent policy $\pi$, message buffer $\mathcal{M}$.
3:for each time_step $t$ in Total_Steps do
4:// — Half-step for agents in $\mathcal{G}_{1}$
5:  Create new empty buffer $\mathcal{M}_{next}$.
6:  for each intersection $i\in\mathcal{G}_{1}$ do
7:   $O_{i},M_{in,i}\leftarrow Env.get\_observation(i),\mathcal{M}[i]$ $\triangleright$ Read messages from $\mathcal{G}_{2}$ from prev. half-step
8:   $A_{i},M_{out,i}\leftarrow\pi(O_{i},M_{in,i})$
9:   if $M_{out,i}\neq\text{null}$ then
10:     for each neighbor $j\in\mathcal{N}(i)\cap\mathcal{G}_{2}$ do
11:      if $dist(i,j)\leq 2\text{km}$ then
12:        $\mathcal{M}_{next}.add(j,M_{out,i})$
13:      end if
14:     end for
15:   end if
16:  end for
17:  $\mathcal{M}\leftarrow\mathcal{M}_{next}$ $\triangleright$ Update message buffer for the next half-step (for $\mathcal{G}_{2}$)
18:  $Env.update\_state(\{A_{i}\}_{i\in\mathcal{G}_{1}})$
19:// — Half-step for agents in $\mathcal{G}_{2}$
20:  Create new empty buffer $\mathcal{M}_{next}$.
21:  for each intersection $j\in\mathcal{G}_{2}$ do
22:   $O_{j},M_{in,j}\leftarrow Env.get\_observation(j),\mathcal{M}[j]$ $\triangleright$ Read messages from $\mathcal{G}_{1}$ from prev. half-step
23:   $A_{j},M_{out,j}\leftarrow\pi(O_{j},M_{in,j})$
24:   if $M_{out,j}\neq\text{null}$ then
25:     for each neighbor $i\in\mathcal{N}(j)\cap\mathcal{G}_{1}$ do
26:      if $dist(j,i)\leq 2\text{km}$ then
27:        $\mathcal{M}_{next}.add(i,M_{out,j})$
28:      end if
29:     end for
30:   end if
31:  end for
32:  $\mathcal{M}\leftarrow\mathcal{M}_{next}$ $\triangleright$ Update message buffer for the next half-step (for $\mathcal{G}_{1}$)
33:  $Env.update\_state(\{A_{j}\}_{j\in\mathcal{G}_{2}})$
34:end for